OpenStudy (anonymous):

what is the limit as x->9 of (9-x)/sqrt(x)-3

OpenStudy (anonymous):

perhaps you can multiply the top and bottom by the conjugate

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

perhaps you are on the right track lagrangeson

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

or we could factor out the top and bottom, yeah that sounds better

OpenStudy (anonymous):

how would i factor the top? there are no exponents

OpenStudy (anonymous):

it's -6

OpenStudy (saifoo.khan):

-_-

OpenStudy (anonymous):

to check just use l'hopitals but 6 is incorrect you're losing a - somewhere

OpenStudy (anonymous):

i made a mistake sorry

OpenStudy (anonymous):

{(9-x)/sqrtx -3} x {sqrtx +3/ sqrt x + 3} {9sqrt x +27-x sqrt x -3x} / {x-9} {- sqrt x (x-9) -3 (x-9)} / {x-9} (x-9) (-sqrt x -3) / (x-9) = -sqrt x -3 = - sqrt 9 -3 = -3-3 =-6

OpenStudy (anonymous):

ok thanks for all your help!

OpenStudy (anonymous):

lagrange had the right idea, there was just a sign mistake.